OpenStudy (anonymous):

Anyone that can help me out on this one please? An investor invested a total of $1,100 in two mutual funds. One fund earned a 7% profit while the other earned a 3% profit. If the investor’s total profit was $61, how much was invested in each mutual fund? The amount that was invested in the mutual fund that earned 7% was? The amount that was invested in the mutual fund that earned 3% was?

OpenStudy (helder_edwin):

u have a linear system to solve: x for the investment in fund 1, y for ... in fund 2 then \[ \large \begin{cases} x+y=1100\\ 0.07x+0.03y=61 \end{cases} \]

OpenStudy (anonymous):

I am a beginner and have no idea what i'm doing, could you explain the steps?

OpenStudy (helder_edwin):

ok. but do u understand the equations? first that!!

OpenStudy (anonymous):

Yes, I see how you have them set up, and I think i'm supposed to start with the first one and isolate x first? but I am unsure of what i am doing as a whole. I would appreciate any help.

OpenStudy (helder_edwin):

well first, to get rid of of the decimal numbers u can multiply the 2nd equation by 100 and get \[ \large 7x+3y=6100 \] ok?

OpenStudy (anonymous):

ok

OpenStudy (helder_edwin):

now isolate x or y (your pick) from the 1st equation. what do u get?

OpenStudy (helder_edwin):

hello! @teenandtoddlermom

OpenStudy (anonymous):

I am here...I am trying to configure it, but I am unsure.

OpenStudy (helder_edwin):

for instance you would get \[ \large x=1100-y \] right?

OpenStudy (anonymous):

yes

OpenStudy (helder_edwin):

now u can plug this into the second equation.and get \[ \large 6100=7x+3y=7(1100-y)+3y \] right?

OpenStudy (anonymous):

yes. I am taking notes on what you are explaining.

OpenStudy (helder_edwin):

can u do the algebra on this last equation: \[ \large 6100=7(1100-y)+3y \] ? post what u get

OpenStudy (anonymous):

ok so ... 7(1100-y)+3 7700-y+3 7700y-3 am I on the right track?

OpenStudy (helder_edwin):

no

OpenStudy (anonymous):

:(

OpenStudy (helder_edwin):

u didn't distribute the 7 properly

OpenStudy (helder_edwin):

let's see: \[ \large 6100=7(1100-y)+3y=7700-7y+3y=7700-4y \] do u see where your mistake is?

OpenStudy (helder_edwin):

so we now have \[ \large 6100-7700=-4y \] \[ \large -1600=-4y \] \[ \large y=\frac{-1600}{-4}=400 \] got it?

OpenStudy (anonymous):

so that is for my first equation? you plugged it into the second equation?

OpenStudy (helder_edwin):

now plug this into the first equation and find the value of x. try it!! it is easy

OpenStudy (anonymous):

so for the x + y =1100 it's 400=1100

OpenStudy (anonymous):

do I subtract them?

OpenStudy (anonymous):

or its, x+400=1100?

OpenStudy (helder_edwin):

what happened with the x?

OpenStudy (helder_edwin):

i mean u now know y=400 so \[ \large 1100=x+y=x+400 \]

OpenStudy (anonymous):

so x=700

OpenStudy (helder_edwin):

yes.

OpenStudy (anonymous):

7% = 700 and 3% is 400 is this right?

OpenStudy (helder_edwin):

no.

OpenStudy (helder_edwin):

x=700 (what u invest in the 1st fund) y=400 (what u invest in the 2nd fund)